OpenStudy (anonymous):

Find an equation of the line that is perpendicular to the line 3x - 5y = 11 and that has the same y-intercept as the line 2x + 9y = 18.

OpenStudy (anonymous):

Rearranging the first eqt.: -5y = 11 - 3x y = (-11/5) + (3/5)x The equation of the perpendicular line has a slope of: -1/M If the slope of the line is (3/5) then the perpendicular line will have a slope: -1/(3/5) = -5/3 therefore the perpendicular equation is: y = (-5/3)x + b We know that the perpendicular line has the same y-intercept as the line 2x + 9y = 18. To find the y-intercept, x=0: 0 + 9y = 18 y = 2 so, y-intercepts at (0,2) Plug in the point into the perpendicular eqt. we find: 2 = 0 + b b = 2 Hence, y = (-5/3)x + 2 Is that right?